OpenStudy (anonymous):

Eliminate the parameter. x=√t, y=2t+5

OpenStudy (jdoe0001):

\(\bf x =\sqrt{t}\qquad y=2t+5 \\ \quad \\ x =\sqrt{t}\implies x^2={\color{red}{ t}}\qquad thus\implies y=2t+5\implies y=2({\color{red}{ x^2}})+5\)

OpenStudy (anonymous):

thanks for the help man :)